Shift report 20030204 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Feb 04 2003 - 00:03:02 EST

  • Next message: Apache: "Shift report 20030204 00:00-08:00"
    ------------------------------------------------------------
    Date: Tue Feb  4 00:03:01 2003, Shift: 16:00-24:00
    Supervisor: kh, with: zm
    Log book#13 pages 146 - 146
    
    Shift Summary:
    ------------------------------------------------------------
    MCR was in the process of dumping the blue beam (deuterons) at the beginning of the shift.  This was for the purpose of doing a background study.  We observed that the TOFW rate dropped by a factor of 5 when the blue beam was dumped.
    
    Plots in the E-Log show the results.
    
    Other than that there was no activity during this shift.
    
    At shift end, both blue and yellow beam had been injected; then:
    "Loss monitor pulled the beam permit link"
    No statement on outlook
    
    Run plan for next shift: (if there is beam)
    MRS:
    Acquire 300K more at 45 deg 400A
    Switch to pol 400B and get 400K
    Move to 40 deg
    400K at 1050B
    400K at 1050A
    Move to 35 deg
    400K at 500A
    400K at 500B
    
    FS:
    Acquire 300K at 1/8B
    Acquire 300K at 1/15B
    Acquire 300K at 1/4B
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Feb 04 2003 - 00:07:19 EST